Objects with masses of 157 kg and 768 kg are separated by 0.342 m. A 28.7 kg mass is placed midway between them. 0.342 m b b 157 kg 28.7 kg 768 kg Find the magnitude of the net gravitational force exerted by the two larger masses on the 28.7 kg mass. The value of the universal gravitational constant is 6.672 × 10−11 N · m2 /kg2 .



1
Expert's answer
2022-01-14T09:47:07-0500

Find the magnitude of the net gravitational force exerted by the two larger masses on the 28.7 kg mass. Below, r is the distance between the 157 kg and 768 kg masses:


"F=|F_1-F_2|=\bigg|\dfrac{4Gm_1m}{r^2}-\dfrac{4Gm_2m}{r^2}\bigg|,\\\space\\
F=\dfrac{4Gm}{r^2}|m_1-m_2|=4·10^{-5}\text{ N}."